# Best-Worst Disaggregation: An approach to the preference disaggregation problem

### v1

@@ Abstract
Preference disaggregation methods in Multi-Criteria Decision-Making (MCDM) often encounter challenges related to inconsistency and cognitive biases when deriving a value function from experts’ holistic preferences. This paper introduces the Best-Worst Disaggregation (BWD) method, a novel approach that integrates the principles of the Best-Worst Method (BWM) into the disaggregation framework to enhance the consistency and reliability of derived preference models. BWD employs the ”consider-the-opposite” strategy from BWM, allowing experts to provide two opposite pairwise comparison vectors of alternatives. This approach reduces cognitive load and mitigates anchoring bias, possibly leading to more reliable criteria weights and attribute value functions. An optimization model is formulated to determine the most suitable additive value function to the preferences expressed by an expert. The method also incorporates a consistency analysis to quantify and improve the reliability of the judgments. Additionally, BWD is extended to handle interval-valued preferences, enhancing its applicability in situations with uncertainty or imprecise information. We also developed an approach to identify a reference set, which is used for pairwise comparisons to elicit the value functions and weights. A case study in logistics performance evaluation demonstrates the practicality and effectiveness of BWD, showing that it produces reliable rankings aligned closely with experts’ preferences.
Keywords: Multi-Criteria Decision-Making (MCDM); Best-Worst Method (BWM); Preference disaggregation; Consistency analysis; Imprecise preference information
@@ 1 Introduction
The field of Multi-Criteria Decision-Making (MCDM) encompasses a wide range of methods designed to aid experts in evaluating a set of alternatives characterized by a set of attributes/criteria (Greco et al., , 2016). Generally, MCDM methods are categorized into two streams: outranking methods and value-based methods (Belton and Stewart, , 2012).
Outranking methods such as ELECTRE (Figueira et al., , 2013) and PROMETHEE (Brans and De Smet, , 2016), compare alternatives pairwise to establish a preference ranking. They determine whether one alternative outranks another based on a set of criteria, focusing on the dominance and credibility of alternatives rather than aggregating criteria scores into a single value.
@@
Value-based methods, such as Multi-Attribute Value Theory (MAVT) (Keeney and Raiffa, , 1993), the Analytic Hierarchy Process (AHP) (Saaty, , 1977), and Best-Worst Method (BWM) (Rezaei, , 2015) aim to assign a numerical value or score to each alternative. In these methods, experts assess the relative importance of criteria and the performance of alternatives on these criteria, often using pairwise comparisons to derive weights and scores. The scores are then aggregated to produce an overall ranking, emphasizing the total utility or value derived from each alternative.
@@
Value based methods are further divided into aggregation and disaggregation approaches. Aggregation methods, such as MAVT, elicit value functions and trade-offs among the objectives to evaluate and rank the alternatives. Disaggregation methods, including the UTA family (Siskos et al., , 2016), on the other hand, are appropriate when an expert can provide holistic preferential information on a subset of alternatives. These methods will then help inferring the underlying value functions and trade-offs. In this context, preference learning emerges as a promising alternative, functioning as an explainable AI that learns and maintains fully interpretable preferences. Preference diaggregation techniques focus on deriving preference structures that are both accurate and transparent, aligning with the growing emphasis on explainable models in decision-making (Rudin, , 2019). By ensuring that the learned preferences are easily understandable, preference learning enhances the usability and acceptance of MCDM methods.
Despite their strengths, disaggregation methods have been subject to critique regarding their consistency and robustness, as they may sometimes struggle to fully address the inconsistencies and biases inherent in human judgment.
@@
Behavioral decision theory suggests that decision-makers often exhibit inconsistencies in their judgment and are influenced by various cognitive biases (Tversky and Kahneman, , 1974). Disaggregation methods, which typically involve only a single round of comparisons, are particularly vulnerable to these inconsistencies. Such limited comparison processes can result in unreliable ranking outcomes due to inconsistent judgments, thereby undermining the reliability of the inferred value functions (Montibeller and Von Winterfeldt, , 2015; Morton and Fasolo, , 2009). This critique is especially relevant for methods like the UTA family, most of which inherently assume a high level of consistency and rationality in the provided preference rankings (Vetschera, , 2006; Schilling et al., , 2007; Korhonen et al., , 2012). While the GRIP (Generalized Regression with Intensities of Preferences) method (Figueira et al., , 2009) represents an exception within the UTA family by considering additional preference information in form of comparisons of intensities of preference between some pairs of reference alternatives, the complexity introduced by integrating regression analysis can make GRIP less transparent and more difficult to implement, especially for practitioners without a strong statistical background.
In addition, a major problem with the practical use of UTA involves incomplete preferences, resulting in multiple or even infinitely many instances of the preference model that align with the expert’s indirect statements (Greco et al., , 2008). This multiplicity can lead to ambiguous recommendations when applied to a set of alternatives. Although preference information can be completed by eliciting additional judgments, this process is often limited in many scenarios (Ciomek et al., , 2017; Salo and Hämäläinen, , 2010). Consequently, robustness analysis methods frequently leaves the problem unsolved, failing to provide a complete ranking or indicate the most preferred alternative. Additionally, analyzing multiple instances of preference models is too abstract for many users, who are more comfortable with a clear, precise solution. To mitigate this, some approaches propose regularization frameworks to balance error minimization with the smoothness of inferred value functions (Doumpos and Zopounidis, , 2007; Liu et al., , 2019). However, these solutions add another layer of complexity, making the methods less accessible to practitioners without advanced technical skills (Kadziński et al., , 2017).
@@
To further address these challenges, this study introduces a novel solution called Best-Worst Disaggregation (BWD) method, which integrates the robustness of BWM within the disaggregation framework of MCDM. The BWD method leverages the structured pairwise comparison approach of BWM alongside the foundational principles of disaggregation methods to derive the value functions, reliably and consistently, from the preferences provided by the experts. The proposed method adopts the consider-the-opposite strategy (Mussweiler and Strack, , 2000; Mussweiler et al., , 2000), by involving comparisons between two opposite reference points—best and worst—and other criteria. This strategy has been incorporated into numerous psychological studies to enhance judgment consistency and demonstrated its efficacy to mitigate the negative effects of the anchoring bias (Adame, , 2016; Joslyn et al., , 2011; Mussweiler, , 2002; Rezaei, , 2021; Rezaei et al., , 2022, 2024). Additionally, the BWD method considers cardinal information, similar to the Deck-of-Cards-Based Ordinal Regression method (DOR) (Corrente et al., 2024b, ). However, while DOR captures both ranking and preference intensity through a combination of Deck of Cards visual elicitation and ordinal regression, allowing it to model complex criteria interactions, its extensive elicitation process can be cumbersome. In contrast, BWD relies on simpler pairwise comparisons, making it more practical and less cognitively demanding for experts.
@@
We summarize the contribution of this study as follows: Introducing a novel preference disaggregation method by combining the Best-Worst Method (BWM) with the disaggregation framework in MCDM, employing the “consider-the-opposite” strategy to reduce cognitive bias, and developing an optimization model to derive the most suitable additive value function based on experts’ pairwise comparisons. Extending the method to handle interval-valued preferences broadens its applicability in situations with uncertainty or imprecise information. A systematic approach is proposed for selecting a reference set that covers all criteria levels, minimizes cognitive load, and avoids Pareto dominance. Consistency analysis is incorporated using both ordinal and cardinal consistency ratios to ensure higher reliability of results, along with sensitivity analysis to assess the maximum and minimum possible ranks for alternatives, evaluating their stability.
@@
The remainder of this paper is structured as follows: Section 2 provides preliminary information on the underlying theories of MCDM, focusing on BWM and UTA methods. Section 3 details the proposed BWD method, illustrating how it integrates the structured pairwise comparison approach of BWM with the disaggregation framework of UTA to derive value functions. In its subsections we explore consistency and sensitivity analysis as well as an interval-valued extensions. Section 4 discusses a real-world application of the method, showcasing its practical utility. Finally, Section 5 concludes the paper and suggests directions for future research.
@@ 2.1 Multi-attribute value theory (MAVT)
Multi-Attribute Value Theory (MAVT) is a framework in MCDM that assumes that each alternative can be characterized by a list of its relevant attributes. A typical MCDM problem consists of a non-empty finite set of $m$ alternatives whose indices range in the set $M=\{1,\ldots,m\}$ and a finite set of $n$ attributes (also known as criteria) indexed by the set $N=\{1,\ldots,n\}$. The set of possible levels that can be achieved by a generic alternative with respect to the $j$th attribute is denoted by $X_{j}$. Assuming all relevant attributes have been considered, each alternative can be associated with a consequence vector $\mathbf{x}=(x_{1},\ldots,x_{n})\in X_{1}\times\cdots\times X_{n}$, where $x_{j}\in X_{j}$ represents the level of the $j$th attribute achieved by the alternative. For simplicity, we will consider consequences vectors $\mathbf{x}$ in place of alternatives. In MAVT, each attribute can be rescaled and normalized into the interval $[0,1]$ using a function $v_{j}:X_{j}\rightarrow[0,1]$ where we assign values 0 and 1 to the least $\underline{x}_{j}$ and most $\overline{x}_{j}$ desirable attribute levels, respectively. Hence, $v_{j}(x_{j})\in[0,1]$ can be interpreted as the value of the consequence evaluated with respect to the $j$th attribute. The primary challenge in MAVT is to find a function $V:[0,1]^{n}\rightarrow[0,1]$ that can correctly aggregate the $n$ values $v_{j}(x_{j})$ and represent the preferences of a decision-maker or an expert. Namely, given a set of alternatives, each of which can be uniquely represented by an associated consequence vector, the value function $V$ should be coherent with the preorder structure on the set of alternatives:
@@
The second family of methods uses elicitation techniques on experts to find compatible attribute value functions $v_{j}$’s and scaling constants $w_{j}$’s. That is, these methods work entirely on the left hand side of equivalence (1). Here we have combinations of methods like the mid-value splitting technique and the trade-off method (Eisenführ et al., , 2010), the SMART (Edwards and Barron, , 1994) and Swing families (Von Winterfeldt and Edwards, , 1986) as well as, among others, the already mentioned AHP and BWM.
@@ 2.2 Best-Worst Method
The Best-Worst Method (BWM) is a pairwise comparison-based method used to elicit the weights of the criteria in a structured way. The pairwise comparison data is collected from the expert who is asked to first identify a set of relevant decision criteria. Then the expert should identify the ‘best’ and the ‘worst’ criteria. The ‘best’ criterion is the one which, according to the expert, has the largest contribution to the goal of the decision-making problem, while the ‘worst’ has the smallest contribution. Then, the expert compares the best criterion, $B$, to all the others, and also all the criteria to the worst one $W$, using a number between 1 to 9 (or possibly some other scales). Finally, an optimization model is formulated and solved to find the weights, $w_{j}$, such that the maximum absolute difference between the pairwise comparisons and their associated weight ratios be minimized. The optimization model is as follows (Rezaei, , 2015),
@@
Solving (4) results in the optimal weights and the optimal objective value can again be used as an inconsistency index. A survey on extensions and applications of the BWM method was proposed by Mi et al., (2019).
@@ 2.3 Preference disaggregation analysis framework
Preference disaggregation is an approach within MCDM that focuses on deriving criteria weights and value functions from some holistic preferences of experts on a reference set of alternatives (Doumpos et al., , 2022). This approach is particularly valuable when:
@@
  • there is a sufficiently large set of well-known alternatives from which an expert or a decision-maker can easily extract a reference set on which she feels sufficiently confident to express her preferences. An example is the study by Corrente et al., (2013) on the innovativeness levels of European countries.
  • comparing some pairs of criteria is difficult. For example, when it comes to choosing a holiday destination, we may be able to easily rank alternatives and yet we may not be able to assess the substitution rate between, for example, distance from home and time taken for the planning.
The UTA (UTilités Additives) method, proposed by Jacquet-Lagreze and Siskos, (1982), is a regression-based approach within the field of MCDM. This method was designed to infer additive value functions that align with the preferences of an expert, as expressed through the ranking of a set of alternatives. The core idea is to disaggregate these preferences to construct value functions that best replicate the given ranking, ensuring consistency across the set of alternatives. The goal of the UTASTAR procedure—an improved variant of the basic UTA method—is to derive additive value functions that reflect the expert’s holistic preferences. First, the expert provides a ranking on a reference set. Next, a linear optimization problem is solved to derive the attribute value functions for each criterion, ensuring that the attribute value functions replicate the provided rankings as closely as possible. More formally, the steps of UTASTAR are the following.
@@
Jacquet-Lagreze and Siskos, (1982) proposed to use piecewise linear approximations of the attribute value functions $v_{j}$. For the $j$th criterion, the interval $[\underline{x}_{j},\overline{x}_{j}]$ can be divided by $s_{j}-1$ equally spaced points into $s_{j}$ sub-intervals, and the approximate attribute value function $v_{j}^{L}$ can be defined as follows. If $x_{j}\in[x_{j}^{k},x_{j}^{k+1}]$, then
@@
$v_{j}^{L}(x_{j}^{k})\leq v_{j}^{L}(x_{j}^{(k+1)})\;\;\forall j\text{ and }\forall k$ (7)
@@
where $x_{j}$ is the attribute level achieved by alternative $\mathbf{x}$ with respect to the $j$th attribute and $\sigma_{\mathbf{x}}^{+}$ and $\sigma_{\mathbf{x}}^{-}$ are overestimation and underestimation errors, respectively.
@@
$\theta=$ minimize $\sum_{i\in R}\sigma_{\mathbf{x}_{i}}^{+}+\sigma_{\mathbf{x}_{i}}^{-}$ (11)
@@
Over the years, several variants and extensions of UTA methods have been developed to enhance its applicability and robustness (Dias et al., , 2018). For instance, the UTADIS variant caters to sorting problems (Zopounidis and Paraschou, , 2013). Another notable extension is the Quasi-UTA model, which simplifies the process by using recursive exponential functions to reduce the information needed to build the utility function (Beuthe et al., , 2000). Since its inception, UTA methods have found applications across various fields (Dias et al., , 2018), including commercial strategy, venture capital investment evaluation, country risk assessment, business financing, and portfolio management (Siskos et al., , 2016). Methods such as UTAGMS and GRIP incorporate robust ordinal regression to handle interacting criteria and enhance the decision-making process (Greco et al., , 2008). These advancements ensure that the UTA methods remain a relevant and powerful tools in the field of MCDM (Doumpos et al., , 2022).

### v2

@@ Abstract
Preference disaggregation methods in Multi-Criteria Decision Analysis (MCDA) often encounter challenges related to inconsistency and cognitive biases when deriving a value function from experts’ holistic preferences. This paper introduces the Best-Worst Disaggregation (BWD) method, a novel approach that integrates the principles of the Best-Worst Method (BWM) into the disaggregation framework to enhance the consistency and reliability of derived preference models. BWD employs the “consider-the-opposite” strategy from BWM, allowing experts to provide two opposite pairwise comparison vectors of alternatives. This approach reduces cognitive load and mitigates anchoring bias, possibly leading to more reliable criteria weights and attribute value functions. An optimization model is formulated to determine the most suitable additive value function to the preferences expressed by an expert. The method also incorporates a consistency analysis to quantify and improve the reliability of the judgments. Additionally, BWD is extended to handle interval-valued preferences, enhancing its applicability in situations with uncertainty or imprecise information. We also developed an approach to identify a reference set, which is used for pairwise comparisons to elicit the value functions and weights. A case study in logistics performance evaluation demonstrates the practicality and effectiveness of BWD, showing that it produces reliable rankings aligned closely with experts’ preferences.
Keywords: Multi-Criteria Decision Analysis (MCDA); Best-Worst Method (BWM); Preference disaggregation; Consistency analysis; Imprecise preference information
@@ 1 Introduction
The field of Multi-Criteria Decision Analysis (MCDA) encompasses a wide range of methods designed to aid experts in evaluating a set of alternatives characterized by a set of attributes/criteria (Greco et al., , 2016). Within MCDA, value-based methods, such as Multi-Attribute Value Theory (MAVT) (Keeney and Raiffa, , 1993), Analytic Hierarchy Process (AHP) (Saaty, , 1977), and Best-Worst Method (BWM) (Rezaei, , 2015) , aim to assign a numerical value to each alternative. In these methods, experts assess the relative importance of criteria and the performance of alternatives on these criteria, often using pairwise comparisons to derive weights and scores. The scores are then aggregated to produce an overall ranking, emphasizing the total utility or value derived from each alternative.
@@
Value-based methods are further divided into aggregation and disaggregation approaches. Aggregation methods, such as MAVT, elicit value functions and trade-offs among the objectives to evaluate and rank the alternatives. Disaggregation methods, including the UTA family (Siskos et al., , 2016), on the other hand, are appropriate when an expert can provide holistic preferential information on a reference set of alternatives. These methods will then help infer the underlying value functions and trade-offs. Doumpos and Zopounidis, (2011) provided a comprehensive review of preference disaggregation methods, emphasizing their connections with statistical learning techniques, and highlighting how these approaches can be integrated to enhance decision support systems. In this context, preference learning emerges as a promising alternative, functioning as an explainable AI that learns and maintains fully interpretable preferences. Preference disaggregation techniques focus on deriving preference structures that are both accurate and transparent, aligning with the growing emphasis on explainable models in decision-making (Rudin, , 2019). By ensuring that the learned preferences are easily understandable, preference learning enhances the usability and acceptance of MCDA methods.
@@
Despite these methodological advances, a fundamental challenge remains. Behavioral decision theory suggests that decision-makers often exhibit inconsistencies in their judgment and are influenced by various cognitive biases (Tversky and Kahneman, , 1974). While disaggregation methods are powerful and widely applicable, their reliance on a single round of holistic comparisons may leave room for judgment inconsistencies to influence the results. The comparison processes might introduce some variability in ranking outcomes, especially when experts’ judgments are internally inconsistent, potentially affecting the stability of the inferred value functions (Montibeller and Von Winterfeldt, , 2015; Morton and Fasolo, , 2009). This consideration is particularly important for methods like the UTA family, most of which inherently assume a high level of consistency and rationality in the provided preference rankings (Vetschera, , 2006; Schilling et al., , 2007; Korhonen et al., , 2012). GRIP (Generalized Regression with Intensities of Preferences) (Figueira et al., , 2009) enriches the UTA family by explicitly modeling preference intensities through a regression framework, thereby offering analysts a more nuanced view of stakeholder judgments; this additional expressive power naturally comes with a need for some statistical expertise and appropriate software support.Barbati et al., (2024) proposed a Deck-of-Cards-based Robust Ordinal Regression (DOR) to push further the use of cardinal preferential information in preference disaggregation methods.
One practical issue that can arise when applying UTA-type methods is that incomplete preference information may leave the analyst with several—occasionally infinitely many—additive value models that all fit the decision maker’s holistic judgments (Greco et al., , 2008). This plurality, while faithful to the data, may translate into more than one plausible recommendation. Eliciting additional judgments is an effective way to narrow the solution set, but time or cognitive-load constraints sometimes limit how much extra information can be gathered (Ciomek et al., , 2017; Salo and Hämäläinen, , 2010). Robust Ordinal Regression and related robustness-analysis procedures offer valuable insight into the full spectrum of admissible rankings, yet presenting a range of models instead of a single solution can feel abstract to users who prefer an immediately actionable outcome. Regularization frameworks—designed to balance goodness-of-fit with smoothness of the inferred value functions—provide another principled way to identify a representative model (Doumpos and Zopounidis, , 2007; Liu et al., , 2019), although their implementation typically demands a higher level of modeling expertise (Kadziński et al., , 2017).
@@
To further address these challenges, this study introduces a novel solution called Best-Worst Disaggregation (BWD) method, which integrates the idea of BWM within the disaggregation framework of MCDA. Let us note that the idea of combining the two approaches was already suggested by Barbati et al., (2024). The BWD method leverages the structured pairwise comparison approach of BWM alongside the foundational principles of disaggregation methods to derive the value functions, reliably and consistently, from the preferences provided by the experts. The proposed method adopts the consider-the-opposite strategy (Mussweiler and Strack, , 2000; Mussweiler et al., , 2000), by involving comparisons between two opposite reference points—best and worst—and other criteria. This strategy has been incorporated into numerous psychological studies to enhance judgment consistency and demonstrated its efficacy to mitigate the negative effects of the anchoring bias (Adame, , 2016; Joslyn et al., , 2011; Mussweiler, , 2002; Rezaei, , 2021; Rezaei et al., , 2022, 2024). While DOR is expressly designed to capture both ranking and preference intensity—thanks to its combination of the Deck-of-Cards visual metaphor and ordinal-regression machinery—this richer elicitation protocol is best suited to decision settings where stakeholders have the time and inclination for in-depth interaction.
@@
We summarize the contribution of this study as follows: Introducing a preference disaggregation method by combining the Best-Worst Method (BWM) with the disaggregation framework in MCDA, employing the “consider-the-opposite” strategy to reduce cognitive bias, and developing an optimization model to derive the most suitable additive value function based on experts’ pairwise comparisons. Extending the method to handle interval-valued preferences broadens its applicability in situations with uncertainty or imprecise information. A systematic approach is proposed for selecting a reference set that covers all criteria levels, minimizes cognitive load, and avoids Pareto dominance. Consistency analysis is incorporated using both ordinal and cardinal consistency ratios to ensure higher reliability of results, along with sensitivity analysis to assess the maximum and minimum possible ranks for alternatives, evaluating their stability.
@@
The remainder of this paper is structured as follows: Section 2 provides preliminary information on the underlying theories of MCDA, focusing on BWM and UTA methods. Section 3 details the proposed BWD method, illustrating how it integrates the structured pairwise comparison approach of BWM with the disaggregation framework of UTA to derive value functions. In its subsections we explore consistency and sensitivity analysis as well as interval-valued extensions. Section 4 discusses a real-world application of the method, showcasing its practical utility. Finally, Section 5 concludes the paper and suggests directions for future research.
@@ 2.1 Multi-attribute value theory (MAVT)
Multi-Attribute Value Theory (MAVT) is a framework in MCDA that assumes that each alternative can be characterized by a list of its relevant attributes. A typical MCDA problem consists of a non-empty finite set of $m$ alternatives whose indices range in the set $M=\{1,\ldots,m\}$ and a finite set of $n$ attributes (also known as criteria) indexed by the set $N=\{1,\ldots,n\}$. The set of possible levels that can be achieved by a generic alternative with respect to the $j$th attribute is denoted by $X_{j}$. Assuming all relevant attributes have been considered, each alternative can be associated with a consequence vector $\mathbf{x}=(x_{1},\ldots,x_{n})\in X_{1}\times\cdots\times X_{n}$, where $x_{j}\in X_{j}$ represents the level of the $j$th attribute achieved by the alternative. For simplicity, we will consider consequence vectors $\mathbf{x}$ in place of alternatives. In MAVT, each attribute can be rescaled and normalized into the interval $[0,1]$ using a function $v_{j}:X_{j}\rightarrow[0,1]$ where we assign values 0 and 1 to the least $\underline{x}_{j}$ and most $\overline{x}_{j}$ desirable attribute levels, respectively. Hence, $v_{j}(x_{j})\in[0,1]$ can be interpreted as the value of the consequence evaluated with respect to the $j$th attribute. The primary challenge in MAVT is to find a function $V:[0,1]^{n}\rightarrow[0,1]$ that can correctly aggregate the $n$ values $v_{j}(x_{j})$ and represent the preferences of a decision-maker or an expert. Namely, given a set of alternatives, each of which can be uniquely represented by an associated consequence vector, the value function $V$ should be coherent with the preorder structure on the set of alternatives:
@@
The second family of methods uses elicitation techniques on experts to find compatible attribute value functions $v_{j}$’s and scaling constants $w_{j}$’s. That is, these methods work entirely on the left-hand side of equivalence (1). Here we have combinations of methods like the mid-value splitting technique and the trade-off method (Eisenführ et al., , 2010), the SMART (Edwards and Barron, , 1994) and Swing families (Von Winterfeldt and Edwards, , 1986) as well as, among others, the already mentioned AHP and BWM.
@@ 2.2 Best-Worst Method
The Best-Worst Method (BWM) is a pairwise comparison-based method used to elicit the weights of the criteria in a structured way. The pairwise comparison data is collected from the expert who is asked to first identify a set of relevant decision criteria. Then the expert should identify the ‘best’ and the ‘worst’ criteria. The ‘best’ criterion is the one which, according to the expert, has the largest contribution to the goal of the decision-making problem, while the ‘worst’ has the smallest contribution. Then, the expert compares the best criterion, $B$, to all the others, and also all the criteria to the worst one $W$, using a number between 1 to 9 (or possibly some other scales). Finally, an optimization model is formulated and solved to find the weights, $w_{j}$, such that the maximum absolute difference between the pairwise comparisons and their associated weight ratios is minimized. The optimization model is as follows (Rezaei, , 2015),
@@
Solving (4) results in the optimal weights and the optimal objective value can again be used as an inconsistency index. A survey on extensions and applications of the BWM method was proposed by Mi et al., (2019). In the BWM, questions on pairs on criteria regard their relative importance in the decision process. Conversely, in MAVT, ratios between weights are interpreted as rates of substitutions. To fill this gap, the Best-Worst Tradeoff (BWT) was introduced (Liang et al., , 2022) to adapt the BWM to the context of MAVT. Whereas the interpretation between the values of the comparisons used in BWM and BWT are different, the formulations of the models are sufficiently similar so that the former is still representative of the latter.
@@ 2.3 Preference disaggregation analysis framework
Preference disaggregation is an approach within MCDA that focuses on deriving criteria weights and value functions from some holistic preferences of experts on a reference set of alternatives (Doumpos et al., , 2022). This approach is particularly valuable when:
@@
  • there is a sufficiently large set of well-known alternatives from which an expert or a decision-maker can easily extract a reference set on which she feels sufficiently confident to express her preferences. An example is the study by Corrente et al., (2013) on the levels of innovation of European countries.
  • comparing some pairs of criteria is difficult. For example, when it comes to choosing a holiday destination, we may be able to easily rank alternatives and yet we may not be able to assess the substitution rate between factors such as distance from home and time taken for the planning.
The UTA (UTilités Additives) method, proposed by Jacquet-Lagreze and Siskos, (1982), is a regression-based approach within the field of MCDA. This method was designed to infer additive value functions that align with the preferences of an expert, as expressed through the ranking of a set of alternatives. The core idea is to disaggregate these preferences to construct value functions that best replicate the given ranking, ensuring consistency across the set of alternatives. The goal of the UTASTAR procedure–an improved variant of the basic UTA method–is to derive additive value functions that reflect the expert’s holistic preferences. First, the expert provides a ranking on a reference set. Next, a linear optimization problem is solved to derive the attribute value functions for each criterion, ensuring that the attribute value functions replicate the provided rankings as closely as possible. More formally, the steps of UTASTAR are the following.
@@
Jacquet-Lagreze and Siskos, (1982) proposed to use piecewise linear approximations of the attribute value functions $v_{j}$. For the $j$th criterion, the interval $[\underline{x}_{j},\overline{x}_{j}]$ can be divided by $(s_{j}-1)$ (with $s_{j}\in\mathbb{N}_{\geq 1}$) equally spaced points in $s_{j}$ subintervals, and the approximate attribute value function $v_{j}^{L}$ can be defined as follows. If $x_{j}\in[x_{j}^{k},x_{j}^{k+1}]$, then
@@
$v_{j}^{L}(x_{j}^{k})\leq v_{j}^{L}(x_{j}^{(k+1)})\;\;\forall j\text{ and }\forall k.$ (7)
@@
where $x_{j}$ is the attribute level achieved by alternative $\mathbf{x}$ with respect to the $j$th attribute and $\sigma_{\mathbf{x}}^{+}$ and $\sigma_{\mathbf{x}}^{-}$ are overestimation and underestimation errors, respectively. Albeit not a strict requirement, pragmatism suggests that the cardinality of $R$ be significantly smaller than the cardinality of $M$.
@@
$\theta=$ minimize $\sum_{i\in R}\left(\sigma_{\mathbf{x}_{i}}^{+}+\sigma_{\mathbf{x}_{i}}^{-}\right)$ (11)
@@
Stability
Given the possible existence of multiple optimal or near-optimal solutions, it is suggested to take the average of the additive value functions that maximize the objective functions $v_{j}(\overline{x}_{j})$ for all $j\in N$, with the same constraints as in (11) and the addition of
$\sum_{i\in R}\left(\sigma_{\mathbf{x}_{i}}^{+}+\sigma_{\mathbf{x}_{i}}^{-}\right)\leq\theta^{*}+\varepsilon$
where $\theta^{*}$ is the optimum of (11) and $\varepsilon>0$ is extremely small, so that only the set of optimal or near-optimal solutions is considered.
Over the years, several variants and extensions of UTA methods have been developed to enhance its applicability and robustness.For instance, the UTADIS variant caters to sorting problems (Zopounidis and Paraschou, , 2013). Another notable extension is the Quasi-UTA model, which simplifies the process by using recursive exponential functions to reduce the information needed to build the utility function (Beuthe et al., , 2000). Since its inception, UTA methods have found applications across various fields (Dias et al., , 2018), including commercial strategy, venture capital investment evaluation, country risk assessment, business financing, and portfolio management (Siskos et al., , 2016). Methods such as UTAGMS and GRIP incorporate robust ordinal regression to handle interacting criteria and enhance the decision-making process (Greco et al., , 2008). These advancements ensure that the UTA methods remain relevant and powerful tools in the field of MCDA (Doumpos et al., , 2022).